You are designing a container in the shape of a cylinder. The radius is 6 inches. You want the container to hold at least 324π cubic inches. What is the least possible height of the container?

Answer by CubeyThePenguin(3113)

V = (pi) * (r^2) * h = 324pi

324pi = pi * (6^2) * h
h = 324/36 = 9

The minimum height is 9 inches.